A piece of aluminum is subjected to a load of 25kn, the aluminum material is cylinderical in shape with a diameter of 42cm. The strain is 25. Calculate the young modulus.
1
Expert's answer
2020-09-06T17:18:45-0400

"\frac{F}{A}=E\epsilon\to E=\frac{F}{A\epsilon}=\frac{4F}{\pi d^2\epsilon}="


"=\frac{4\cdot 25000}{3.14\cdot 0.42^2\cdot25}=7221(Pa)"